2(p + 5) - (9 + p) = -3. I can not figure out how the book got the answer of -4. Could you please show all the steps invloved?
Thanks,
Brandy

Answer by Abbey(339)

Since I don't know where you made your error, I can guess that either you forgot to distribute the negative sign in front of the (9+p)or you forgot to distribute the 2 to the (p+5), so:
using the distributive property we will clear the parentheses:
2(p + 5) - (9 + p) = -3
2p+10-9-p=-3
then combine like terms:
p+1=-3
subtract 1 from both sides:
p=-4